# LLM-Coordination: Evaluating and Analyzing Multi-agent Coordination Abilities in Large Language Models

### v2

@@ Abstract
The emergent reasoning and Theory of Mind (ToM) abilities demonstrated by Large Language Models (LLMs) make them promising candidates for developing coordination agents. In this study, we introduce a new LLM-Coordination Benchmark aimed at a detailed analysis of LLMs within the context of Pure Coordination Games, where participating agents need to cooperate for the most gain. This benchmark evaluates LLMs through two distinct tasks: (1) Agentic Coordination, where LLMs act as proactive participants for cooperation in 4 pure coordination games; (2) Coordination Question Answering (QA), where LLMs are prompted to answer 198 multiple-choice questions from the 4 games for evaluation of three key reasoning abilities: Environment Comprehension, ToM Reasoning, and Joint Planning. Furthermore, to enable LLMs for multi-agent coordination, we introduce a Cognitive Architecture for Coordination (CAC) framework that can easily integrate different LLMs as plug-and-play modules for pure coordination games. Our findings indicate that LLM agents equipped with GPT-4-turbo achieve comparable performance to state-of-the-art reinforcement learning methods in games that require commonsense actions based on the environment. Besides, zero-shot coordination experiments reveal that, unlike RL methods, LLM agents are robust to new unseen partners. However, results on Coordination QA show a large room for improvement in the Theory of Mind reasoning and joint planning abilities of LLMs. The analysis also sheds light on how the ability of LLMs to understand their environment and their partner’s beliefs and intentions plays a part in their ability to plan for coordination. Our code is available at https://github.com/eric-ai-lab/llm_coordination.
@@ 1 Introduction
In a wide range of activities, from daily tasks such as cooking to critical operations like rescue efforts, cooperation without mixed intentions is essential. These scenarios are examples of Pure Coordination Games, where all involved parties benefit from choosing strategies that are perfectly aligned, avoiding any conflict of interest. These games require agents to reason about their environment and plan while considering the beliefs and intentions of their partners. Recently, Large Language Models (LLMs) have demonstrated emergent planning abilities in both physical and virtual settings (Raman et al., 2022; Wang et al., 2023; Wu et al., 2023), impressive reasoning capabilities (Wei et al., 2022), and the hints of a Theory of Mind (Kosinski, 2023) making them promising candidates for developing coordination agents. Previous works have explored the use of LLMs for developing collaborative agents, yet the requisite conditions, strengths, and limitations of LLMs in coordination games remain unclear. In this study, we intend to bridge the gap by performing a comprehensive evaluation and analysis of the multi-agent coordination abilities of LLMs.
@@
Therefore, we introduce a new LLM-Coordination Benchmark featuring two task settings for pure coordination games: 1. Agentic Coordination and 2. CoordinationQA. In Agentic Coordination, LLMs are scaffolded with components that allow them to act within actual game environments, providing a holistic evaluation of the competencies of LLMs to act as coordination agents. In CoordinationQA, LLMs have to answer a curated set of questions about edge-case scenarios drawn from coordination games where agents are required actively cooperate with their partners. The benchmark includes four collaborative games, providing a comprehensive analysis platform.
To enable LLMs for multi-agent coordination, we present a Cognitive Architecture for Coordination (CAC) framework that facilitates LLM interaction with game environments in a plug-and-play approach. CAC translates game elements into textual formats and leverages auxiliary LLMs for improved coordination to enable effective multi-agent collaboration. Our Experiments on the Agentic Coordination task with CAC reveal that Large Language Models are capable of understanding the game objectives, generating coherent reasoning for their next actions, and coordinating with partners across all coordination games. They exhibit these competencies without any training, fine-tuning, or few-shot examples.
@@
A comparative analysis on the Agentic Coordination task reveals that LLM agents outperform Multi-agent RL solutions in games that require minimum theory-of-mind reasoning and focus on taking commonsense actions based on the environment. However, they struggle in more complex settings where agents need to actively consider their partner’s beliefs and intentions. We also observe that LLM agents are capable of collaborating with new partners, unlike self-play MARL methods (Carroll et al., 2019a; Bard et al., 2020) that fail to adapt to unseen agents.
@@
For a more nuanced analysis of the coordination abilities of LLMs, we create the CoordinationQA Suite. This suite is designed to dissect the capabilities of LLMs in single-turn reasoning within coordination games, focusing on three key areas: Joint Planning, Theory of Mind (ToM), and Environment Comprehension. Joint Planning evaluates LLMs’ decision-making for optimal long-term outcomes, ToM questions probe the understanding of partner agents’ intentions and needs, and Environment Comprehension assesses knowledge of game rules and dynamics. Our findings on CoordinationQA show a marked performance gap between GPT-4-turbo and other LLMs across three question types. LLMs are most proficient in Environment Comprehension, indicating they understand game rules and states well. However, they face significant challenges in Theory of Mind Reasoning, with difficulty inferring others’ intentions and needs. This issue worsens in Joint Planning, where most LLMs underperform, some even worse than random choices. These results highlight LLMs’ limited reliability and effectiveness as coordination partners. Additionally, our analysis shows a moderate to strong correlation between Theory of Mind Reasoning, Environment Comprehension, and Joint Planning performance, underscoring their role in effective coordination.
In summary, our contributions are four-fold:
@@
We develop the plug-and-play Cognitive Architecture for Coordination Framework to enable LLMs to effectively participate in complex, partially observable coordination games like Hanabi, demonstrating the first zero-shot application for LLM agents in such scenarios.
@@
4.
We investigate Environment Comprehension and Theory of Mind Reasoning as essential components of LLMs’ overall Joint Planning capabilities, highlighting their critical importance in coordination tasks.
@@ 2.1 Multi-turn Agentic Coordination
In the Multi-turn Agentic Coordination task, LLMs participate in end-to-end pure coordination games as agents, where the best strategy for all participating agents is to cooperate. LLMs under test are plugged into coordination frameworks with components like memory modules and grounding modules to act in complete games. These LLM agents can then be partnered with any policies or agents to complete the games.
Our LLM-Coordination benchmark includes 4 pure coordination games, Hanabi Challenge (Bard et al., 2020), Overcooked-AI (Carroll et al., 2019a), Collab Capture and Collab Escape.
@@
Hanabi Challenge: In Hanabi (Bard et al., 2020), players aim to assemble five sequences of cards in ascending order (1 through 5). A unique aspect of the game is that the players can only view their partner’s cards, not their own. This requires players to work collaboratively, utilizing reveal tokens to provide hints about the cards in their partner’s hand. These hints can be about either the color or the rank of the cards. For instance, using a single hint token, a player can indicate all cards of a certain rank in their partner’s hand. Hanabi serves as an exemplary Pure Coordination game, necessitating player cooperation to achieve optimal outcomes. Success in Hanabi hinges on the ability to understand partners’ perspectives, navigate decisions based on incomplete information, and engage in implicit communication, making it an excellent testing ground for coordination among agents.
@@
Overcooked-AI: In the Overcooked-AI environment (Carroll et al., 2019a), two agents—Alice (Blue) and Bob (Green)—collaborate to cook and deliver onion soups. This environment includes a variety of layouts, each with its own arrangement and quantity of onion dispensers, plate dispensers, cookers, delivery zones, and countertops. To prepare a dish, agents are required to insert three onions into a cooker, initiating a cooking process that lasts 20 time steps. Upon completion, the soup must be plated and delivered to complete the task. Each layout presents unique challenges, emphasizing the need for agents to comprehend their surroundings, locate necessary resources, and synchronize their actions with their teammate for effective collaboration.
Collab Capture: Collab Capture involves two agents trying to capture an adversary in a maze of interconnected rooms. The rooms are connected by doors, which can be controlled through access buttons that can be found in other rooms. The agents’ task is to capture the adversary in the least amount of time using effective strategies, including cornering the adversary and manipulating the doors to enable their partner or disable the adversary.
@@
Collab Escape: Collab Escape involves two agents trying to escape an adversary in a maze of interconnected rooms. They need to fix two generators (similar to the game Dead-by-Daylight (Dea, 2016)) located in different rooms to open an exit portal. The adversary tries to catch the agents, and the win condition is any one agent escaping. This game requires strategies like luring the adversary away from the partner, sacrificing for the partner’s safety, and manipulating the movement of the adversary.
@@ 2.2 Single-turn Coordination QA
Environment Comprehension (EC) questions require LLMs to make indirect inferences about some aspect of their environment.
Theory of Mind Reasoning (ToM) questions challenge the LLMs to predict the intentions of their partners and probe about the requirements of their partners.
Joint Planning (JP) questions provide agents with the state/observation and ask them to predict the best next action for effective coordination. This question is essentially the same question that LLMs need to repeatedly solve when they act as agents.
All the questions were manually developed and labeled. We filtered out questions and scenarios that showed any ambiguity, leaving only questions that had clear optimal solutions. We generated a total of N=66 scenarios (25 from Overcooked, 28 from Hanabi, and 13 from the two Collab Games) and created 3 questions per scenario, resulting in 198 unique questions. The right side of Figure 1 demonstrates the sampling process for the three types of questions with an example from the game Overcooked. The selected scenario shows the Blue agent about to place their third onion in the cooker, and the green agent needs to figure out what to do next.
3 Cognitive Architecture for Coordination
We develop a LLM Agent architecture based on Sumers et al. (2023), which we dub Cognitive Architecture for Coordination (CAC), for multi-agent coordination. Using CAC we can easily plug and play a LLM agent, and pair it with any partner (e.g., another CAC agent, a human player, or other AI agents). The architecture consists of three key elements: Memory, Reasoning, and Grounding.
Memory Module includes (1) Long-Term Memory for storing the Game Description including the game’s rules, conventions, objectives and action space, (2) Working memory which consists a textual description of the current observation, and (3) Episodic Memory which is a list of previous actions selected by the agent. In the example shown in Figure 2 the Long Term Memory includes a description of the game of Hanabi, The Working memory includes observations including the current stack, partner’s hand, beliefs and knowledge of both agents and information regarding the available tokens, and cards, and the Episodic Memory includes the previously selected discards, plays and hints.
Reasoning Module is where the Large Language Model (LLM) is plugged into the framework. It takes the textual description in the working memory as input and generates the next best action based on the context. For coordination games like Hanabi that require a more sophisticated Theory of Mind reasoning, we add an auxiliary Theory of Mind Reasoning LLM whose sole responsibility is to interpret the partner agent’s actions and requirements. This extra reasoning is added to the working memory before passing to the primary LLM. Additionally, we also utilize a Self-verification LLM which verifies the safety of the selected action. In the Hanabi example in Figure 2, the reasoning module interprets the provided clue of ”Revealing Rank 1 Cards” and decides to play one of the pointed cards on the empty stacks. The generated action is passed to the grounding module.
Grounding Module is responsible for interfacing the reasoning and memory modules’ textual decision-making spaces with the actual game mechanics. Its primary task is to take the selected action from the reasoning module and translate it into game-compatible action(s). The exact implementation of the grounding module depends on the game in question; for example, in Overcooked-AI, the grounding module needs to convert high-level actions like ”pick up onion from o0.” into sequences of lower-level actions. On the other hand, in games like Hanabi, it just needs to match actions like ”Reveal Bob’s Red Color Cards” to their lower-level representations. The grounding module is also responsible for the secondary task of filtering out infeasible actions based on the context of the game.
4 Experiments

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated emergent common-sense reasoning and Theory of Mind (ToM) capabilities, making them promising candidates for developing coordination agents. This study introduces the LLM-Coordination Benchmark, a novel benchmark for analyzing LLMs in the context of Pure Coordination Settings, where agents must cooperate to maximize gains. Our benchmark evaluates LLMs through two distinct tasks. The first is Agentic Coordination, where LLMs act as proactive participants in four pure coordination games. The second is Coordination Question Answering (CoordQA), which tests LLMs on 198 multiple-choice questions across these games to evaluate three key abilities: Environment Comprehension, ToM Reasoning, and Joint Planning. Results from Agentic Coordination experiments reveal that LLM-Agents excel in multi-agent coordination settings where decision-making primarily relies on environmental variables but face challenges in scenarios requiring active consideration of partners’ beliefs and intentions. The CoordQA experiments further highlight significant room for improvement in LLMs’ Theory of Mind reasoning and joint planning capabilities. Zero-Shot Coordination (ZSC) experiments in the Agentic Coordination setting demonstrate that LLM agents, unlike RL methods, exhibit robustness to unseen partners. These findings indicate the potential of LLMs as Agents in pure coordination setups and underscore areas for improvement. Code Available at https://github.com/eric-ai-lab/llm_coordination.
LLM-Coordination: Evaluating and Analyzing Multi-agent Coordination Abilities in Large Language Models
@@ 1 Introduction
In a wide range of activities, from daily tasks such as cooking to critical operations like rescue efforts, cooperation without mixed intentions is essential. These scenarios are examples of Pure Coordination Games, where all involved parties benefit from choosing strategies that are perfectly aligned, avoiding any conflict of interest. These games require agents to reason about their environment and plan while considering the beliefs and intentions of their partners. Recently, Large Language Models (LLMs) have demonstrated emergent planning abilities in both physical and virtual settings (Raman et al., 2022; Wang et al., 2023a; Wu et al., 2023), impressive reasoning capabilities (Wei et al., 2022), and the hints of a Theory of Mind (Kosinski, 2023) making them promising candidates for developing coordination agents. Previous works have explored the use of LLMs for developing collaborative agents, yet the requisite conditions, strengths, and limitations of LLMs in coordination games remain unclear. In this study, we intend to bridge the gap by performing a comprehensive evaluation and analysis of the multi-agent coordination abilities of LLMs.
@@
Therefore, we introduce the LLM-Coordination Benchmark featuring two task settings for pure coordination games: 1. Agentic Coordination and 2. CoordinationQA. In Agentic Coordination, LLMs are scaffolded with components that allow them to act within actual game environments, providing a holistic evaluation of the competencies of LLMs to act as coordination agents. In CoordinationQA, LLMs have to answer a curated set of questions about edge-case scenarios drawn from coordination games where agents need to actively cooperate with their partners. The benchmark includes four collaborative games, providing a comprehensive analysis platform. Unlike studies on multi-LLM frameworks Hong et al. (2023); Qian et al. (2024); Li et al. (2023a), which focus on orchestrating multiple LLMs to solve tasks, our benchmark assesses the innate ability of individual LLMs to understand and act within pure coordination scenarios where cooperation is essential.
@@
Our experiments in the Agentic Coordination setting reveal that Large Language Models are competent at understanding the game objectives, generating coherent reasoning for their next actions, and coordinating with partners across all coordination games. They exhibit these behaviors without any training, fine-tuning, or few-shot examples. A comparative analysis reveals that LLM agents match or outperform RL baselines in games where optimal decision-making can be done by observing environment variables and positions (e.g., Overcooked). However, they struggle in settings where agents need to actively consider their partner’s beliefs and intentions (e.g., Hanabi). We also observe that LLM agents are capable of collaborating with new partners, unlike self-play MARL methods (Carroll et al., 2019a; Bard et al., 2020) that fail to adapt to unseen agents.
@@
For a more nuanced analysis of the coordination abilities of LLMs, we create the CoordinationQA Suite. This suite is designed to dissect the capabilities of LLMs in single-turn reasoning within coordination games, focusing on three key areas: Joint Planning, Theory of Mind (ToM), and Environment Comprehension. Joint Planning (JP) evaluates LLMs’ planning abilities for optimal coordination, ToM questions probe their understanding of partner agents’ intentions and needs, and Environment Comprehension (EC) assesses their inference of environment details, rules and objectives. First, Our findings on CoordinationQA show a marked performance gap between GPT-4-turbo and other LLMs across three question types. Secondly, LLMs are most proficient in Environment Comprehension, indicating they understand the rules and environment states well. However, they face significant challenges in Theory of Mind Reasoning, with difficulty inferring others’ intentions and needs. This issue worsens in Joint Planning, where most LLMs underperform, some even worse than random choices. These results highlight LLMs’ limited reliability and effectiveness as coordination partners. Correlation analysis between LLMs’ performance on CoordinationQA and their performance on agentic coordination setting further highlights their strengths in environmental reasoning but exposes significant weaknesses in Theory of Mind inference and Joint Planning capabilities.
In summary, our contributions are threefold:
@@
We investigate Environment Comprehension, Theory of Mind Reasoning, and Joint Planning as essential components of LLMs’ overall coordination capabilities, highlighting their critical importance in pure coordination setups.
2 Related Work
Multi-agent Coordination.
Pure Coordination games in game theory are scenarios where agents share the payoffs, and cooperation is the optimal strategy. Benchmarks like the Multiparticle Environment Lowe et al. (2017), Overcooked-AI Carroll et al. (2019a), and the Hanabi Challenge Bard et al. (2020) evaluate multi-agent coordination. Carroll et al. (2019a) highlighted the importance of human data for effective Human-AI collaboration. Subsequent Overcooked-AI research focuses on aligning self-play-trained agents with humans using techniques such as self-play with past checkpoints (Strouse et al., 2021), population entropy objectives (Zhao et al., 2023), graph-theoretic objectives (Li et al., 2023c), policy ensembles (Lou et al., 2023), and integrating human biases (Yu et al., 2023). In the Hanabi Challenge, efforts aim to learn grounded policies over arbitrary conventions Hu et al. (2021b, a). While most solutions enhance RL methods for coordination, we propose that LLMs offer an alternative due to their emergent reasoning and theory-of-mind-like abilities, avoiding arbitrary joint interactions.
Planning and Reasoning with Large Language Models.
LLMs have shown remarkable natural language reasoning abilities (OpenAI, 2023; Ouyang et al., 2022; Chiang et al., 2023), achieving state-of-the-art results in verbal reasoning tasks. Augmented with components like memory and tools, LLMs can interact with external environments, solving long-horizon tasks and playing complex games (Wu et al., 2023; Wang et al., 2023a; Liang et al., 2022; Song et al., 2022). Guided by Cognitive Architectures for Language Agents (Sumers et al., 2023) as a design principle for agent design, we experiment with advanced reasoning strategies such as ReAct Yao et al. (2023), Self-Verification Weng et al. (2023), and Self-Consistency Wang et al. (2023b) to enhance LLM reasoning. These strategies establish strong baseline performance for our Language Agent implementations.
Multi-agent LLMs.
Recent studies have explored LLMs in multi-agent cooperation settings. Zhang et al. (2023b) developed a modular agent framework for spatial rearrangement tasks. Zhang et al. (2023a) introduced an architecture enabling LLMs to play Overcooked-AI. Shi et al. (2023) demonstrated positive zero-shot coordination in Avalon using code-driven reasoning. Li et al. (2023d) showed emergent collaborative abilities of LLMs in simulations, while Li et al. (2023b) investigated theory-of-mind inference using explicit belief representations. Xu et al. (2024) perform an analysis of LLMs in communication games Xu et al. (2023) analyzed LLM cognitive abilities through games, highlighting benefits of probabilistic modeling. In contrast, our research rigorously evaluates LLM agents’ coordination abilities in established pure coordination games, where coordination is essential. Our setting only includes scenarios where agents must fully cooperate with each other with no competitive incentives. We also conduct a fine-grained, component-level analysis to understand the intricacies of LLMs’ coordination capabilities.
@@ 3.1 Multi-turn Agentic Coordination
In the Multi-turn Agentic Coordination task, LLMs participate in end-to-end pure coordination games as agents, where the best strategy for all participating agents is to cooperate. We only consider pure coordination scenarios with no competitive incentives. LLMs under test are plugged into coordination frameworks with memory and the ability to act in complete games. These LLM agents can then be partnered with any policies or agents to complete the games.
Our LLM-Coordination benchmark includes 4 pure coordination games: Hanabi Challenge (Bard et al., 2020), Overcooked-AI (Carroll et al., 2019a), and Collab Capture and Collab Escape (inspired by the Pursuit-Evasion problem). These games were carefully selected for their ability to isolate and highlight specific coordination challenges, providing controlled environments that allow the analysis of pure coordination scenarios without too much emphasis on other reasoning challenges. While LLMs are versatile and capable of addressing a wide range of tasks, these well-studied settings offer established benchmarks, clear metrics, and reproducible scenarios that are particularly suited for examining coordination abilities of participating agents.
Hanabi Challenge.
@@
In Hanabi (Bard et al., 2020), players aim to assemble five sequences of cards in ascending order (1 through 5), each sequence dedicated to a different color: purple, red, blue, yellow, and green. A unique aspect of the game is that the players can only view their partner’s cards, not their own. This requires players to work collaboratively, utilizing reveal tokens to provide hints about the cards in their partner’s hand. These hints can be about either the color or the rank of the cards. For instance, using a single reveal token, a player can indicate all cards of a certain rank in their partner’s hand. Once a player has an idea about which card they have, they can choose to play the card on the stack. If the card is correct, they get a point. Otherwise, players lose a collective life token. The loss of all 3 life tokens leads to the end of the game. Hanabi serves as an exemplary Pure Coordination game, necessitating player cooperation to achieve optimal outcomes. Success in Hanabi hinges on the ability to understand partners’ perspectives, navigate decisions based on incomplete information, and engage in implicit communication, making it an excellent testing ground for coordination among agents.
Overcooked-AI.
@@
In the Overcooked-AI environment (Carroll et al., 2019a), two agents—Alice (Blue) and Bob (Green)—collaborate to cook and deliver onion soups. This environment includes a variety of layouts, each with its own arrangement and quantity of onion dispensers, plate dispensers, cookers, delivery zones, and countertops. To prepare a dish, agents are required to insert three onions into a cooker, initiating a cooking process that lasts 20 time steps. Upon completion, the soup must be plated and delivered to complete the task. Each layout presents unique challenges, emphasizing the need for agents to comprehend their surroundings, locate necessary resources, and synchronize their actions with their teammate for effective collaboration.
Collab Capture.
Collab Capture involves two agents trying to capture an adversary in a maze of interconnected rooms. The rooms are connected by doors, which can be controlled through access buttons that can be found in other rooms. The agents’ task is to capture the adversary in the least amount of time using effective strategies. To evaluate different coordination strategies between agents, we design four scenarios by controlling the states of the doors (open or closed) within the layout shown in figure 5. These scenarios highlight various coordination challenges, such as trapping the adversary through precise positioning, enabling a teammate by prioritizing door control over direct pursuit, and strategically restricting the adversary’s movement to facilitate capture. (see Appendix E.1 for more details.)
Collab Escape.
@@
Collab Escape involves two agents trying to escape an adversary in a maze of interconnected rooms. They need to fix two generators (similar to the game Dead-by-Daylight (Dea, 2016)) located in different rooms to open an exit portal. The adversary tries to catch the agents, and the win condition is any one agent escaping. To evaluate coordination strategies in Collab Escape, we develop two scenarios by varying the initial proximity of agents to the adversary and generators in the layout shown in figure 6. Depending on their proximity to the adversary/generators, players need to apply strategies such as luring the adversary away from the partner, choosing to continue fixing the generators while sacrificing for the partner’s safety and manipulating the movement of the adversary (see Appendix E.2 for more details.)
@@ 3.2 Single-turn Coordination QA
Environment Comprehension (EC) questions require LLMs to make indirect inferences about some aspect of their environment (See Appendix D.1). The questions cover details of the layouts, implications of current observations, and counts of artifacts.
Theory of Mind Reasoning (ToM) questions challenge the LLMs to predict the intentions of their partners and probe about the requirements of their partners (See Appendix D.2).
Joint Planning (JP) questions provide agents with the state/observation and ask them to predict the best next action for effective coordination. This question is essentially the same question that LLMs need to repeatedly solve when they act as agents (See Appendix D.3).
All the questions were manually developed and labeled. We filtered out questions and scenarios that showed any ambiguity, leaving only questions that had clear, optimal solutions. We generated a total of N=66 scenarios (25 from Overcooked, 28 from Hanabi, and 13 from the two Collab Games) and created 3 questions per scenario, resulting in 198 unique questions. The right side of Figure 1 demonstrates the sampling process for the three types of questions with an example from the game Overcooked. The selected scenario shows the Blue agent about to place their third onion in the cooker, and the green agent needs to figure out what to do next. See Appendix D for examples of questions and the templates used to formulate these questions.
4 Experimental Setup
@@ 4.1 Agentic Coordination
4.1.1 LLM Agents
To allow LLMs to play multi-turn games, we scaffold them with an agentic framework based on Cognitive Architectures for Language Agents Sumers et al. (2023). The framework includes three parts: Memory, Reasoning, and Grounding.
Memory includes (1) Long-Term Memory for storing the Game Description, including the game’s rules, conventions, objectives, and action space, (2) Working memory, which consists of a textual description of the current observation, and (3) Episodic Memory which is a list of previous actions selected by the agent.
Reasoning is where the Large Language Model (LLM) is plugged into the framework. It takes the textual description from the Memory as input and generates the next action based on the context. The LLM reasons about the current state and then selects an action from the list of available actions in natural language.
Self-Verification: For the coordination game Hanabi, there is a low margin for error as any misplays lead to the loss of life tokens, and the loss of all three life tokens subsequently results at the end of the game. We thus supplement the reasoning process in Hanabi with Answer-Verification Weng et al. (2023), where the LLM is re-prompted to confirm that the action it generated is appropriate and does not lead to fatal errors.
ToM-Reasoning: We also demonstrate the positive impact of a Theory of Mind Reasoning step prior to generating the next action for Hanabi and CollabEscape, which benefit from this intermediate step. In the ToM reasoning step, the LLM generates an interpretation of their partner’s actions or current position before generating the next action to explicitly capture the belief inference process. We do not test with additional ToM reasoning on Overcooked due to significant latency and cost constraints, with marginal benefits.
Finally, the Grounding process translates the natural language action generated by the reasoning module into game-compatible action(s). The exact implementation of the grounding module depends on the game in question; for example, in Overcooked-AI, the grounding module needs to convert high-level actions like "pick up onion from o0." into sequences of lower-level actions. On the other hand, in games like Hanabi, the Grounding needs to match actions like "Reveal Bob’s Red Color Cards" to their lower-level representations. The Grounding process is also responsible for filtering out infeasible actions based on the context of the game (See Appendices A, B for more details.)
There are no prompt or setup differences for LLM Agents based on Cross-play or Self-play. We use the LLMs gpt-4-0125-preview, GPT-3.5-turbo-0125, Mixtral 8x7B, and GPT-4o for agentic evaluation studies.
| Agent | CR | AA | Ring | FC | CC |
| Self-Play (PPO) | |||||
| PBT | |||||
| GPT-3.5-turbo | |||||
| Mixtral8x7B | |||||
| GPT-4o | |||||
| GPT-4-turbo |
| Collab Escape | Collab Capture | |||
| Agent | Capture Rate | Avg. Turns | Escape Rate | Avg. Turns |
| GPT-4-turbo | 0.83 | 4.60 | 1.00 | 3.5 |
| – (w/out ToM Reasoning) | 0.50 | 2.00 | 1.00 | 4.75 |
| GPT-4o | 0.67 | 4.00 | 1.00 | 7.17 |
| GPT-3.5-turbo | 0.33 | 2.50 | 0.67 | 8.38 |
| Mixtral-8x7b | 0.50 | 7.67 | 0.92 | 7.55 |
| Greedy Baseline | 0.00 | N.A. | 0.50 | 6.00 |
4.1.2 MARL Agents